# A Conflicts-free, Speed-lossless KAN-based Reinforcement Learning Decision System for Interactive Driving in Roundabouts

### v1

@@ VI-C Hard Mode Validation
In this validation, we rigorously assess the proposed system’s driving safety and efficiency within the most challenging operational context. Our methodology contrasts the system’s performance against benchmark algorithms through a two-phase validation process. The first validation focuses on the system’s ability to achieve rapid convergence during training, maintain stability under diverse speed variations, and minimize collision occurrences. The second validation analyzes reward metrics during the evaluation period, highlighting the system’s capacity for optimal decision-making and overall effectiveness during evaluation. Collectively, these validations are designed to substantiate the proposed system’s advanced learning efficiency and robust safety features, positioning it as a superior solution in interactive roundabout driving.
Validation 1: Better training performance. We employed three distinct random seeds to generate a variety of scenarios, ensuring a robust evaluation against other benchmark algorithms. The results, depicted in Fig. 10, reveal a clear superiority of the K-DQN in terms of training performance. Specifically, Fig. 10(a) demonstrates that the K-DQN not only achieved significantly higher peak rewards—approximately 160, compared to the nearest competitor’s 100—but also exhibited a markedly faster convergence rate. In stark contrast, the conventional DQN lagged with rewards near 75. Fig. 10(b) highlights the K-DQN’s proficiency in maintaining both higher and more stable speeds during training, peaking at around 23 m/s, whereas the PPO struggled with the lowest and most erratic speeds, oscillating between 11 m/s and 16 m/s. The A2C, ACKTR, and DQN displayed comparable trajectories, with the A2C slightly edging out with speeds up to 18 m/s after 10,000 training episodes. Lastly, Fig. 10(c) underscores the K-DQN’s enhanced safety feature, registering a collision rate below 0.05, a significant reduction compared to the DQN’s 0.52, ACKTR’s 0.35, and the more variable rates of A2C and PPO, which hovered around 0.2 and 0.15, respectively. These findings collectively affirm the K-DQN’s exceptional training efficacy and potential for real-world application due to its high rewards, speed stability, and low collision rates.
Figure 12: Curves of reward during evaluation of K-DQN, and benchmark algorithms on hard mode scenario.
@@
Validation 2: Reward values among evaluation. In the conducted evaluation, the proposed K-DQN algorithm’s reward efficacy was benchmarked against conventional algorithms using a triad of random seeds to foster diverse scenario generation, as illustrated in Fig. 12. The empirical data unequivocally demonstrates the K-DQN’s superior reward acquisition, consistently outperforming its counterparts. The traditional DQN algorithm trailed with suboptimal rewards, not exceeding 70. In contrast, both the A2C and ACKTR algorithms exhibited marginal improvements, hovering around a reward threshold of 75. The PPO algorithm, while achieving a moderate reward peak of 125, generally maintained a reward level around 100. Remarkably, the K-DQN’s reward performance remained robust, with a notable fluctuation around 175, thereby establishing its dominance over the other evaluated benchmark algorithms in terms of reward maximization.
Table III provides a comparative analysis of the testing performance for collision rate and average speed between the proposed K-DQN method and four SOTA benchmarks: PPO, A2C, ACKTR, and DQN, evaluated in normal mode and Hard Mode scenarios. In normal mode, the proposed K-DQN demonstrates the lowest collision rate at 0.01, significantly outperforming PPO at 0.23, A2C at 0.21, ACKTR at 0.28, and DQN at 0.52, which indicates the least safe performance by DQN. Additionally, the proposed method records the highest average speed of 21.59 m/s, indicating superior efficiency compared to PPO at 14.76 m/s, A2C at 18.83 m/s, ACKTR at 17.89 m/s, and DQN at 18.31 m/s. In Hard Mode, the proposed K-DQN again shows the lowest collision rate at 0.02, maintaining a clear advantage in safety over PPO at 0.12, A2C at 0.19, ACKTR at 0.31, and DQN at 0.52. It also achieves the highest average speed of 22.52 m/s, demonstrating excellent performance in more challenging scenarios, compared to PPO at 13.67 m/s, A2C at 18.04 m/s, ACKTR at 17.53 m/s, and DQN at 17.70 m/s. The proposed K-DQN consistently shows the lowest collision rate across both modes, underscoring its superior safety mechanisms, and achieves the highest average speed, indicating greater efficiency and stability compared to the benchmark algorithms. The proposed K-DQN consistently shows the lowest collision rate across both modes, underscoring its superior safety mechanisms, and achieves the highest average speed, indicating greater efficiency and stability compared to the benchmark algorithms.
In this study, we have used the average speed of vehicles as a key performance indicator, which serves as an indirect measure of the roundabout capacity. Higher average speeds generally indicate a more efficient flow of traffic and, consequently, a higher capacity. By comparing the average speed of our proposed K-DQN method with benchmark algorithms, we have demonstrated the effectiveness of our approach in improving the overall efficiency and capacity of the roundabout. Among the benchmarks, PPO shows moderate performance but falls behind the proposed method in both collision rate and average speed, while A2C and ACKTR perform better than PPO and DQN but do not reach the performance level of the proposed K-DQN. DQN exhibits the poorest performance in terms of collision rate, though it maintains a relatively high average speed. Overall, the proposed K-DQN method demonstrates significant improvements in both safety, as indicated by lower collision rates, and efficiency, as indicated by higher average speeds, compared to the four benchmark algorithms in both easy and hard driving scenarios, making it highly effective for safe and efficient driving in complex environments.
@@ VII Conclusion
This paper presents a DRL-based algorithm to enhance the safety and efficiency of AVs navigating roundabouts, especially in complex traffic scenarios involving HDVs. By utilizing a DQN capable of processing state information of surrounding vehicles, the system eliminates the need for manual feature engineering and thus enables the AVs to effectively interpret their environment. The integration of a KAN significantly bolsters the AVs’ capacity to learn their surroundings with greater accuracy and reliability. The algorithm also incorporates an action inspector to minimize collision risks and a route planner to improve driving efficiency. Additionally, the MPC control ensures the stability and precision of driving actions, contributing to the overall robustness of the system. Evaluations demonstrate the algorithm’s superior performance, consistently achieving safe and efficient driving across varying traffic flows. Compared to state-of-the-art benchmark algorithms, the proposed algorithm shows a notable reduction in collision incidents, decreased travel times, and accelerated training convergence rates. Future research will aim to: 1) test our algorithm’s robustness in more unpredictable traffic scenarios, including complex urban ramps and intersections, 2) enhance the driving strategy by integrating further goals such as improved passenger comfort, and 3) expand the algorithm’s applicability to include collaborative multi-agent operations, which will help in understanding its performance in team-based and competitive urban traffic situations.

### v2

@@ VI-C Hard Mode Validation
This validation assesses the proposed system’s safety and efficiency in the most challenging conditions, with eleven initial vehicles in the roundabout as depicted in Fig. 5(b).
Figure 11: Rewards for K-DQN and benchmarks: hard mode.
[…]
Validation 2: Reward values among evaluation. The evaluation, using three random seeds to ensure scenario diversity (Fig. 11), highlights the K-DQN algorithm’s superior reward performance. Traditional DQN remained below 70, while A2C and ACKTR show slight improvements near 75. PPO reaches a peak of 125 but averaged around 100. In contrast, K-DQN consistently outperforms all baselines, maintaining reward levels around 175, demonstrating clear dominance in reward maximization.
@@
Table III compares collision rates and average speeds of the proposed K-DQN against PPO, A2C, ACKTR, and DQN in both normal and hard modes. In normal mode, K-DQN achieves the lowest collision rate (0.01), outperforming PPO (0.23), A2C (0.21), ACKTR (0.28), and DQN (0.52). It also records the highest average speed at $21.59~\mathrm{m/s}$, surpassing PPO ($14.76~\mathrm{m/s}$), A2C ($18.83~\mathrm{m/s}$), ACKTR ($17.89~\mathrm{m/s}$), and DQN ($18.31~\mathrm{m/s}$). In hard mode, K-DQN maintains its advantage with the lowest collision rate (0.02) and highest speed ($22.52~\mathrm{m/s}$), while all benchmarks show reduced safety and efficiency. These results confirm K-DQN’s superior safety, efficiency, and robustness across varying traffic complexities.
This study employs average vehicle speed as a key performance metric, reflecting roundabout capacity and traffic efficiency. Higher average speeds imply improved flow and increased capacity. Comparative analysis shows that the proposed K-DQN consistently outperforms benchmark algorithms in both safety and efficiency. While PPO exhibits moderate performance, it lags behind K-DQN in both metrics. A2C and ACKTR perform better than PPO and DQN but fall short of K-DQN. DQN records the highest collision rate, despite maintaining relatively high speeds. Overall, K-DQN achieves lower collision rates and higher average speeds across both normal and hard scenarios, highlighting its effectiveness for safe, efficient navigation in complex traffic environments.
@@ VII Conclusion
This paper proposes a DRL-based algorithm to improve AV safety and efficiency in complex roundabout traffic with HDVs. Using a DQN that processes surrounding vehicle states, it avoids manual feature engineering and enhances environmental perception. The integration of a KAN further improves learning accuracy and reliability. The algorithm includes an action inspector to reduce collisions, a route planner for efficient driving, and MPC control for stable, precise actions. Evaluations show superior performance with fewer collisions, reduced travel times, and faster training convergence compared to state-of-the-art benchmarks. Future research will focus on: 1) evaluating the algorithm’s robustness in more complex traffic scenarios, including urban ramps with elevation changes, multi-lane intersections with pedestrian crossings, and roundabouts with varied lane widths ($3.5\text{--}5.0~\mathrm{m}$) and radii to assess adaptability to diverse infrastructures; 2) enhancing driving strategies by incorporating passenger comfort (jerk minimization) and energy efficiency for electric vehicles; and 3) extending the algorithm to collaborative multi-agent settings with V2V communication and platooning to evaluate performance in cooperative and competitive urban traffic.